%Working 68 45 73 54 61
%Management 49 21 66 47 33

Develop the estimated regression equation by computing the values b0 and b1.

Enter negative values as negative numbers. Round your answers to one decimal place.
ŷ = ? + ? x

d. Predict the percentage of management jobs held by women in a company that has 60% women employees. Round your answers to nearest whole value.
? %
Answer #1

)b) b0=-35.0 , b1=1.3 and

(c) ŷ = b0 + b1*x= - 35.0+1.3*x

(d) 43% of management jobs held by women in a company that has 60% women employees.

for x=60, ŷ=-35.0+1.3*60=43%

%Working(x) %Management(y)
68 49
45 21
73 66
54 47
61 33
mean= 60.2 43.2
variance 98.96 232.96
covariance 128.56
b1=cov(x,y)/var(x)= 1.299110752
b0=mean(y)-b1*mean(x)= -35.0065